# Confidence on the Focal: Conformal Prediction with Selection-Conditional Coverage

### v1

@@ B.7 Proof of Proposition 6
    where the first step is because $\mathds{1}\{\widehat{S}_{i}\geq T_{j}^{\textnormal{fixed}}(1,0),Y_{i}\leq c_{i}\}=1$ and $y\leq c_{n+j}$; the second step is due to the definition of $T_{j}^{\textnormal{fixed}}(1,0)$. Eqn. (B.7) implies that $T_{j}^{\textnormal{fixed}}(1,0)\geq T^{{\textnormal{swap}(i,j)}}(y)$. Conversely, since $\mathds{1}\{\widehat{S}_{i}\geq T^{{\textnormal{swap}(i,j)}}(y),Y_{i}\leq c_{i}\}=1$, we have
    $\sum_{\ell\in{\mathcal{I}}_{\textnormal{calib}}}\mathds{1}\{\widehat{S}_{\ell}\geq T^{\textnormal{swap}(i,j)}(y),Y_{\ell}\leq c_{\ell}\}+\mathds{1}\{\widehat{S}_{n+j}\geq T^{\textnormal{swap}(i,j)}(y)\}$
[…]
    following the definition of $T^{\textnormal{swap}(i,j)}(y)$. We then conclude that $T^{\textnormal{swap}(i,j)}(y)\geq T^{\textnormal{fixed}}_{j}(1,0)$ and therefore $T_{j}^{(1,0)}=T^{{\textnormal{swap}(i,j)}}(y)$. As a result,
    $\{\ell\in[m]:\widehat{S}_{n+\ell}^{{\textnormal{swap}(i,j)}}\geq T_{j}^{\textnormal{fixed}}(1,0)\}=\{\ell\in[m]:\widehat{S}_{n+\ell}^{{\textnormal{swap}(i,j)}}\geq T^{\textnormal{swap}(i,j)}(y)\}=\widehat{{\mathcal{S}}}^{{\textnormal{swap}(i,j)}}(y)\in\mathfrak{S}.$
    We have therefore verified that $i\in\widehat{{\mathcal{R}}}^{\textnormal{fixed}}_{n+j}(y)$.
    Next, given $i\in\widehat{{\mathcal{R}}}^{\textnormal{fixed}}_{n+j}(y)$, our goal now is to show that $i\in\widehat{{\mathcal{R}}}_{n+j}(y)$. By definition, there is
    $\widehat{S}_{i}\geq T_{j}^{\textnormal{fixed}}(1,0)\text{ and }\{\ell\in[m]:\widehat{S}_{n+\ell}^{\textnormal{swap}(i,j)}\geq T_{j}^{\textnormal{fixed}}(1,0)\}\in\mathfrak{S}.$
    Since $y\leq c_{n+j}$ and $\mathds{1}\{\widehat{S}_{i}\geq T^{\textnormal{fixed}}_{j}(1,0),Y_{i}\leq c_{i}\}=1$,
    $1+\sum_{\ell\in{\mathcal{I}}_{\textnormal{calib}}\backslash\{i\}}\mathds{1}\{\widehat{S}_{\ell}\geq T_{j}^{\textnormal{fixed}}(1,0),Y_{\ell}\leq c_{\ell}\}+\mathds{1}\{\widehat{{\mathcal{S}}}_{n+j}\geq T_{j}^{\textnormal{fixed}}(1,0),y\leq c_{n+j}\}$
[…]
    where the last step is due to the definition of $T_{j}^{\textnormal{fixed}}(1,0)$. The above implies that $T_{j}^{\textnormal{fixed}}(1,0)\geq T^{\textnormal{swap}(i,j)}(y)$, and therefore $\widehat{S}_{i}\geq T^{\textnormal{swap}(i,j)}(y)$. Conversely, since $\mathds{1}\{\widehat{S}_{i}\geq T^{\textnormal{swap}(i,j)}(y),Y_{i}\leq c_{i}\}=1$, and $y\leq c_{n+j}$,
    $\sum_{\ell\in{\mathcal{I}}_{\textnormal{calib}}}\mathds{1}\{\widehat{S}_{\ell}\geq T^{\textnormal{swap}(i,j)}(y),Y_{\ell}\leq c_{\ell}\}+\mathds{1}\{\widehat{S}_{n+j}\geq T_{j}^{\textnormal{swap}(i,j)}(y)\}$
@@
    following the definition of $T^{\textnormal{swap}(i,j)}(y)$. As a result, $T^{\textnormal{swap}(i,j)}(y)\geq T_{j}^{\textnormal{fixed}}(1,0)\Rightarrow T_{j}^{\textnormal{fixed}}(1,0)=T^{\textnormal{swap}(i,j)}(y)$. The swapped selection set
    $\widehat{{\mathcal{S}}}^{\textnormal{swap}(i,j)}(y)$ =$=\{\ell\in[m]:\widehat{S}_{n+\ell}^{\textnormal{swap}(i,j)}\geq T^{\textnormal{swap}(i,j)}(y)\}=\{\ell\in[m]:\widehat{S}_{n+\ell}^{\textnormal{swap}(i,j)}\geq T_{j}^{\textnormal{fixed}}(1,0)\}\in\mathfrak{S}.$
    It follows that $i\in\widehat{{\mathcal{R}}}_{n+j}(y)$.
  • $Y_{i}>c_{i}$.
    Note that for any $t\in{\mathcal{T}}$,
    $1+\sum_{\ell\in{\mathcal{I}}_{\textnormal{calib}}}\mathds{1}\{\widehat{S}_{\ell}\geq t,Y_{\ell}\leq c_{\ell}\}+\mathds{1}\{\widehat{S}_{n+j}\geq t\}$
[…]
    becuase $y\leq c_{n+j}$ and $Y_{i}>c_{i}$. Inspecting the definition of $T_{j}^{\textnormal{fixed}}(1,1)$ and $T^{\textnormal{swap}(i,j)}(y)$, we can conclude that $T_{j}^{\textnormal{fixed}}(1,1)=T^{\textnormal{swap}(i,j)}(y)$. As a result, we have $\{\widehat{S}_{i}\geq T^{\textnormal{fixed}}_{j}(1,1)\}=\{\widehat{S}_{i}\geq T^{\textnormal{swap}(i,j)}(y)\}$ and
    $\{\ell\in[m]:\widehat{S}_{\ell}^{\textnormal{swap}(i,j)}\geq T_{j}^{\textnormal{fixed}}(1,1)\}=\{\ell\in[m]:\widehat{S}_{\ell}^{\textnormal{swap}(i,j)}\geq T_{j}^{\textnormal{swap}(i,j)}(y)\}.$
    Therefore $i\in\widehat{{\mathcal{R}}}_{n+y}(y)\iff i\in\widehat{{\mathcal{R}}}_{n+j}^{\textnormal{fixed}}(y)$.
Combing the cases of $Y_{i}\leq c_{i}$ and $Y_{i}>c_{i}$, we have shown that ^$\widehat{{\mathcal{R}}}_{n+j}(y)=\widehat{{\mathcal{R}}}^{\textnormal{fixed}}_{n+j}(y)$ for $y\leq c_{n+j}$.
Case 2: $y>c_{n+j}$.
We again consider the case of $Y_{i}\leq c_{i}$ and $Y_{i}>c_{i}$ separately.
  • $Y_{i}\leq c_{i}$.
    Fix $i\in\widehat{{\mathcal{R}}}_{n+j}(y)$. By definition, there is $\widehat{S}_{i}\geq T^{\textnormal{swap}(i,j)}(y)$ and $\widehat{{\mathcal{S}}}^{\textnormal{swap}(i,j)}(y)\in\mathfrak{S}$. Since $y>c_{n+j}$,
    $\sum_{\ell\in{\mathcal{I}}_{\textnormal{calib}}}\mathds{1}\{\widehat{S}_{\ell}\geq T^{\textnormal{swap}(i,j)}(y),Y_{\ell}\leq c_{\ell}\}$
@@
    where the last step follows from the definition of $T^{\textnormal{swap}(i,j)}(y)$. We thus have $T^{\textnormal{swap}(i,j)}(y)\geq T_{j}^{\textnormal{fixed}}(0,0)$ and $\widehat{S}_{i}\geq T_{j}^{\textnormal{fixed}}(0,0)$. Conversely, since $\mathds{1}\{\widehat{S}_{i}\geq T_{j}^{\textnormal{fixed}}(0,0),Y_{i}\leq c_{i}\}=1$,
    $1+\sum_{\ell\in{\mathcal{I}}_{\textnormal{calib}}\backslash\{i\}}\mathds{1}\{\widehat{S}_{\ell}\geq T_{j}^{\textnormal{fixed}}(0,0),Y_{\ell}\leq c_{\ell}\}+\mathds{1}\{\widehat{S}_{n+j}\geq T_{j}^{\textnormal{fixed}}(0,0),y\leq c_{n+j}\}$
@@
    by the definition of $T_{j}^{\textnormal{fixed}}(0,0)$. As a result, we have $T_{j}^{\textnormal{fixed}}(0,0)=T^{\textnormal{swap}(i,j)}(y)$, and
    $\{\ell\in[m]:\widehat{S}_{n+\ell}^{{\textnormal{swap}(i,j)}}\geq T_{j}^{\textnormal{fixed}}(0,0)\}=\widehat{{\mathcal{S}}}^{\textnormal{swap}(i,j)}(y)\in\mathfrak{S}.$
    It then follows that $i\in\widehat{{\mathcal{R}}}^{\textnormal{fixed}}_{n+j}(y)$.
    Conversely, if $i\in\widehat{{\mathcal{R}}}^{\textnormal{fixed}}_{n+j}(y)$, then $\widehat{S}_{i}\geq T_{j}^{\textnormal{fixed}}(0,0)$ and $\{\ell\in[m]:\widehat{S}_{n+\ell}^{\textnormal{swap}(i,j)}\geq T_{j}^{\textnormal{fixed}}(0,0)\}\in\mathfrak{S}.$ Note that since $\mathds{1}\{\widehat{S}_{i}\geq T_{j}^{\textnormal{fixed}}(0,0),Y_{i}\leq c_{i}\}=1$ and $y>c_{n+j}$
    $1+\sum_{\ell\in{\mathcal{I}}_{\textnormal{calib}}\backslash\{i\}}\mathds{1}\{\widehat{S}_{\ell}\geq T_{j}^{\textnormal{fixed}}(0,0),Y_{\ell}\leq c_{\ell}\}+\mathds{1}\{\widehat{S}_{n+j}\geq T_{j}^{\textnormal{fixed}}(0,0),y\leq c_{n+j}\}$
@@
    following the definition of $T_{j}^{\textnormal{fixed}}(0,0)$. As a result, we have $T_{j}^{\textnormal{fixed}}(0,0)\geq T^{\textnormal{swap}(i,j)}(y)$. Conversely, since $\mathds{1}\{\widehat{S}_{i}\geq T^{\textnormal{swap}(i,j)}(y),Y_{i}\leq c_{i}\}=1$,
    $\sum_{\ell\in{\mathcal{I}}_{\textnormal{calib}}}\mathds{1}\{\widehat{S}_{\ell}\geq T^{\textnormal{swap}(i,j)}(y),Y_{\ell}\leq c_{\ell}\}$

### v3

@@ E.1 Connection between selection-conditional coverage and FCR control
That is, the asymptotic FCR coincides with the asymptotic selection-conditional miscoverage.
Proof.
By the definition of FCR, we have
FCR $=\mathbb{E}\bigg{[}\frac{\sum_{j=1}^{m}\mathds{1}\{j\in\widehat{\mathcal{S}},V_{n+j}>\widehat{q}_{m,n}\}}{1\vee|\widehat{\mathcal{S}}|}\bigg{]}$
$=\mathbb{E}\bigg{[}\frac{\sum_{j=1}^{m}\mathds{1}\{j\in\widehat{\mathcal{S}}^{*},V_{n+j}>\widehat{q}_{m,n}\}}{1\vee|\widehat{\mathcal{S}}|}\bigg{]}$
$\qquad-\mathbb{E}\bigg{[}\frac{\sum_{j=1}^{m}\mathds{1}\{j\in\widehat{\mathcal{S}}^{*}\setminus\widehat{\mathcal{S}},V_{n+j}>\widehat{q}_{m,n}\}}{1\vee|\widehat{\mathcal{S}}|}\bigg{]}+\mathbb{E}\bigg{[}\frac{\sum_{j=1}^{m}\mathds{1}\{j\in\widehat{\mathcal{S}}\setminus\widehat{\mathcal{S}}^{*},V_{n+j}>\widehat{q}_{m,n}\}}{1\vee|\widehat{\mathcal{S}}|}\bigg{]}.$ (47)
Note that
$\Bigg{|}-\mathbb{E}\bigg{[}\frac{\sum_{j=1}^{m}\mathds{1}\{j\in\widehat{\mathcal{S}}^{*}\setminus\widehat{\mathcal{S}},V_{n+j}>\widehat{q}_{m,n}\}}{1\vee|\widehat{\mathcal{S}}|}\Bigg{]}+\mathbb{E}\bigg{[}\frac{\sum_{j=1}^{m}\mathds{1}\{j\in\widehat{\mathcal{S}}\setminus\widehat{\mathcal{S}}^{*},V_{n+j}>\widehat{q}_{m,n}\}}{1\vee|\widehat{\mathcal{S}}|}\bigg{]}\bigg{|}$
@@
which goes to zero by assumption and the dominated convergence theorem. For the first term in Equation (E.1), we further have
$\mathbb{E}\bigg{[}\frac{\sum_{j=1}^{m}\mathds{1}\{j\in\widehat{\mathcal{S}}^{*},V_{n+j}>\widehat{q}_{m,n}\}}{1\vee|\widehat{\mathcal{S}}|}\bigg{]}$
@@
The absolute value of the second term above is bounded by $\mathbb{E}[|\widehat{\mathcal{S}}^{*}\Delta\widehat{\mathcal{S}}|/|\widehat{\mathcal{S}}|]$, and therefore goes to zero by assumption. Next, by the strong law of large numbers,
@@
Meanwhile,
$\bigg{|}\frac{1}{m}\sum_{j=1}^{m}\mathds{1}\{j\in\widehat{\mathcal{S}}^{*},V_{n+j}>\widehat{q}_{m,n}\}-G(\widehat{q}_{m,n})\bigg{|}$ $\leq\sup_{t\in\mathbb{R}}\leavevmode\nobreak\ \bigg{|}\frac{1}{m}\sum_{j=1}^{m}\mathds{1}\{j\in\widehat{\mathcal{S}}^{*},V_{n+j}>t\}-G(t)\bigg{|}$
@@
The last step follows from the uniform law of large numbers. Since $G(t)$ is continuous in $t$ and that $\widehat{q}_{m,n}\stackrel{{\scriptstyle p}}{{\to}}q^{*}$, there is
$\frac{1}{m}\sum_{j=1}^{m}\mathds{1}\{j\in\widehat{\mathcal{S}}^{*},V_{n+j}>\widehat{q}_{m,n}\}\stackrel{{\scriptstyle\text{p}}}{{\to}}G(q^{*}).$ (48)
Then, applying the continuous mapping theorem and the dominated convergence theorem yields
$\mathbb{E}\bigg{[}\frac{\sum_{j=1}^{m}\mathds{1}\{j\in\widehat{\mathcal{S}}^{*},V_{n+j}>\widehat{q}_{m,n}\}}{1\vee|\widehat{\mathcal{S}}^{*}|}\bigg{]}\rightarrow\mathbb{P}\big{(}V(X_{n+1},Y_{n+1})>q^{*}{\,|\,}f(X_{n+1})\geq\tau\big{)}.$
This establishes the expression of the asymptotic FCR. In addition, due to the exchangeability of $\{Z_{n+j}\}_{j=1}^{m}$ and that $\widehat{S}$ is equivariant to the permutation of the test points, we have